then A If A=matrix A-A 2 alues eigenua are its #s? imaginary
Answer #1

solution; Geiren that A = matrix and Let C = A-AT 2 Now T CT T-((^-?) CT = (A-AT) 2 1 :: A+B)T=te! (QAFQAT (AT)T-A AGIR CT